### v1

Abstract
The main purpose of this work is to present a SIMD-vectorized implementation of the symplectic 16th-order 8-stage implicit Runge-Kutta integrator based on collocation with Gauss-Legendre nodes (IRKGL16-SIMD), and to show that it can outperform state-of-the-art symplectic explicit integrators for high-precision numerical integrations (in double-precision floating-point arithmetic) of non-stiff Hamiltonian ODE systems. Our IRKGL16-SIMD integrator leverages Single Instruction Multiple Data (SIMD) based parallelism (in a way that is transparent to the user) to significantly enhance the performance of the sequential IRKGL16 implementation. We present numerical experiments comparing IRKGL16-SIMD with state-of-the-art high-order explicit symplectic methods for the numerical integration of several Hamiltonian systems in double-precision floating-point arithmetic.
keywords:
Symplectic methods, Gauss implicit Runge-Kutta methods, SIMD-vectorization, non-stiff Hamiltonian ODE systems, Julia implementation
1 Introduction
The family of $s$-stage implicit Runge-Kutta schemes based on collocation with Gauss-Legendre nodes (IRKGL) are known to be symmetric, symplectic and super-convergent of order $2s$, and hence are very well suited for high precision long-term numerical integration of Hamiltonian systems [18, 11]. The computational cost per step of an implicit RK method is typically higher than that of an explicit RK method, as a system of algebraic equations must be solved at each step of implicit integrators. For non-stiff problems, fixed-point iterations are commonly employed to approximately solve such systems of algebraic equations at each step. A fixed-point implementation (based on the ideas presented in [1] and written in Julia language) of the $16$th order $8$-stage IRKGL scheme (IRKGL16) is included in the DifferentialEquations.jl suite [2, 16]. Our code, IRKGL16, includes a standard time-step adaptivity mechanism. However, for long-term integration of Hamiltonian systems, the use of a constant time step is recommended [18, 11, 13]. Therefore, in this work, we will focus on its application in constant step-size mode.
While there are no explicit symplectic Runge-Kutta schemes, optimized high-order explicit symplectic integrators are available for Hamiltonian systems with suitable structure. These integrators, that can be derived with the techniques of splitting and composition, are applicable when the Hamiltonian function can be decomposed into the sum of two, three, or more exactly solvable Hamiltonians [5, 11, 13].
It is widely acknowledged that high-order explicit symplectic methods are more efficient than IRKGL schemes for computations in double-precision floating-point arithmetic [11]. In fact, to achieve similar accuracy over a given integration interval, our sequential implementation of IRKGL16 usually requires more CPU time than the most effective high order explicit symplectic integrators in double-precision computations.
However, it is well established [11] that the efficiency of implicit Runge-Kutta methods can be significantly enhanced through various forms of parallelism. Single Instruction Multiple Data (SIMD), often referred to as SIMD-vectorization, is a type of parallelism that can be leveraged in most modern CPU cores, enabling vector computations at a cost comparable to scalar operations. The structure of the IRKGL16 integrator is particularly well-suited for exploiting SIMD parallelism with vectors containing eight (or alternatively, four or two) floating-point numbers.
The primary objective of this work is to present a SIMD-vectorized implementation of IRKGL16 in constant step-size mode (referred to as IRKGL16-SIMD) and to demonstrate its superiority over state-of-the-art explicit symplectic integrators for high-precision numerical integrations in double-precision floating-point arithmetic for several relevant Hamiltonian problems with a separable structure.
Our implementation makes use of the Julia package SIMD [19] to explicitly handle computations on vectors containing a fixed number of floating-point numbers. The algorithm IRKGL16-SIMD automatically performs vectorization, ensuring that the process remains transparent to the user.
The integrator IRKGL16-SIMD is designed to numerically integrate non-stiff systems of (non-necessarily Hamiltonian) ODEs whose right-hand side is implemented as a generic Julia function using arithmetic operations and elementary functions supported by the SIMD package.
The paper is organized as follows: Section 1 provides an introduction; Section 2 describes a vectorized fixed-point implementation of symplectic implicit Runge-Kutta (IRK) schemes; Section 3 presents implementation details of the IRKGL16-SIMD method; Section 4 reviews state-of-the-art symplectic explicit integrators; Section 5 reports numerical experiments; and Section 6 presents the conclusions.
2 Vectorized fixed-point implementation of symplectic IRK schemes
2.1 Implicit RK methods
We consider initial value problems for systems of ordinary differential equations (ODEs) of the form:
$\frac{d}{dt}y=f(t,y),\quad y(t_{0})=y_{0},$ (1)
where $f:\mathbb{R}^{D+1}\to\mathbb{R}^{D}$ is a sufficiently smooth function and $y_{0}\in\mathbb{R}^{D}$.
Runge-Kutta (RK) methods are widely used one-step integrators for numerically solving ODEs. These methods are specified by the coefficients in the Butcher tableau:
[…]
The RK approximations $y_{n}\approx y(t_{n})$ to the solution $y(t)$ at times $t_{n}=t_{0}+nh$ for $n=1,2,3,\dots$ are given by:
$y_{n}=y_{n-1}+h\sum_{i=1}^{s}b_{i}f(t_{n-1}+c_{i}h,Y_{n,i}),$ (2)
where
$Y_{n,i}=y_{n-1}+h\sum_{j=1}^{s}a_{ij}f(t_{n-1}+c_{j}h,Y_{n,j}),\quad i=1,\ldots,s.$ (3)
Here, each $Y_{n,i}$ is an approximation of the state vector $y(t)$ at the intermediate time $t=t_{n-1}+hc_{i}$ computed within each time-step. The coefficients $a_{ij}$, $b_{i}$, and $c_{i}$ from the Butcher tableau determine the specific RK method and its properties.
If the matrix $A=(a_{ij})$ is lower triangular the Runge-Kutta method is explicit. Otherwise, it is an implicit Runge-Kutta (IRK) scheme, and the stage vectors $Y_{n,i}$ at each step are defined implicitly by the equation (3).
2.2 IRK methods of collocation type
In a collocation method based on the nodes $c_{1},\ldots,c_{s}\in[0,1]$, the approximations $y_{n}\approx y(t_{n})$ to the solution $y(t)$ at times $t_{n}=t_{0}+nh$ are obtained as follows: for each $n=1,2,3,\dots$, consider the polynomial function $P_{n}:\mathbb{R}\to\mathbb{R}^{D}$ of degree $s$ uniquely determined by the following conditions,
  • $P_{n}(t_{n-1})=y_{n-1}$,
  • $P_{n}^{\prime}(t_{n-1}+c_{i}h)=f(t_{n-1}+c_{i}h,P_{n}(t_{n-1}+c_{i}h))$ for $i=1,\ldots,s$.
then, set $y_{n}=P_{n}(t_{n})$.
This is equivalent [11] to the application of the IRK method (2)–(3) with coefficients $b_{i},a_{ij}$ uniquely determined by the following conditions:
$\begin{split}\sum_{j=1}^{s}a_{ij}c_{j}^{k-1}&=\frac{c_{i}^{k}}{k},\quad i=1,\ldots,s,\quad k=1,2,\ldots,s,\\ \sum_{j=1}^{s}b_{j}c_{j}^{k-1}&=\frac{1}{k},\quad k=1,2,\ldots,s.\end{split}$ (4)
In the specific case where $c_{1},\ldots,c_{s}\in(0,1)$ are the shifted Gauss-Legendre nodes–defined by $c_{i}=(1+x_{i})/2$, where $x_{1},\ldots,x_{s}$ are the zeroes of the Legendre polynomial of degree ss–the resulting IRK scheme of collocation type, which we refer to as the $s$-stage IRKGL scheme, achieves an order of accuracy $2s$ [11].
2.3 Symplectic IRK methods with floating-point coefficients
We consider an autonomous Hamiltonian system of the form
$\frac{dy}{dt}=J^{-1}\nabla H(y),\quad y\in\mathbb{R}^{2d},$ (5)
where $H:\mathbb{R}^{2d}\to\mathbb{R}$ is the Hamiltonian function and $J$ is the standard symplectic matrix
$J=\begin{pmatrix}0&I_{d}\\ -I_{d}&0\end{pmatrix}$.
The exact flow of this system, denoted by $\varphi_{H}^{t}$, is symplectic, meaning that it preserves the symplectic form: for all $t$, the Jacobian matrix $D\varphi_{H}^{t}(y)$ satisfies
$\left(D\varphi_{H}^{t}(y)\right)^{T}JD\varphi_{H}^{t}(y)=J.$
It is therefore desirable that numerical integrators used to approximate such flows also preserve this structure, in order to ensure long-time stability and accurate qualitative behavior of the solution.
An implicit Runge-Kutta (IRK) method is symplectic if and only if [18] its coefficients satisfy the condition
$b_{i}a_{ij}+b_{j}a_{ji}=b_{i}b_{j},\quad 1\leqslant i,j\leqslant s.$ (6)
This condition ensures that the discrete flow defined by the IRK method is symplectic, and is satisfied in particular by the $s$-stage IRKGL schemes [18, 13].
As shown in [12], if the coefficients $b_{i},a_{ij}$ are replaced by machine numbers $\tilde{b}_{i},\tilde{a}_{ij}$, then the resulting implicit RK scheme typically fails to satisfy the simplectic condition (6). To ensure that the symplectic property is exactly preserved with double-precision floating-point coefficients, we reformulate the IRKGL schemes as follows [1]:
$Y_{n,i}$ $=y_{n-1}+\sum^{s}_{j=1}{\mu_{ij}\,L_{n,j},\quad L_{n,i}=hb_{i}f(t_{n-1}+hc_{i},Y_{n,i})},\quad i=1,\ldots,s,$ (7)
$y_{n}$ $=y_{n-1}+\sum^{s}_{i=1}L_{n,i},$ (8)
where $\mu_{ij}=a_{ij}/b_{j},\quad 1\leqslant i,j\leqslant s$. If the coefficients $b_{i}$, $c_{i}$, and $\mu_{ij}$ are replaced by the double-precision floating-point coefficients $\tilde{b}_{i}$, $\tilde{c}_{i}$, and $\tilde{\mu}_{ij}$, then the resulting scheme is still symplectic provided that $\tilde{\mu}_{ij}$ is set for $1\leqslant i<j\leqslant s$ as $1-\tilde{\mu}_{ji}$.
2.4 Vectorized fixed-point iteration
The primary challenge in implementing implicit Runge-Kutta (IRK) methods is the efficient solution of the nonlinear system for the internal stages. For non-stiff problems, fixed-point iteration is generally recommended. In collocation-type IRK methods, it is standard practice to initialize the fixed-point iteration with an estimate of the internal stages extrapolated from the previous time-step, that is,
$Y_{n,i}\approx P_{n-1}(t_{n-1}+hc_{i}),\quad i=1,\ldots,s.$ (9)
It is not difficult to check that
$P_{n-1}(t_{n-1}+hc_{i})=y_{n-1}+\sum_{j=1}^{s}\nu_{ij}L_{n-1,j},\quad i=1,\ldots,s,$ (10)
where the coefficients $\nu_{ij}$ are uniquely determined from the following equations
$\sum_{j=1}^{s}\nu_{ij}(c_{i}-1)^{k-1}=\frac{c_{i}^{k}}{k},\quad i=1,\ldots,s,\quad k=1,\ldots,s.$
Vectorized notation
We introduce notation to facilitate the description of the proposed vectorized implementation of the fixed-point iteration.
Let $b$ and $c$ denote the $s$-vectors $(b_{1},\ldots,b_{s})$ and $(c_{1},\ldots,c_{s})$, respectively. Similarly, for each $i=1,\ldots,s$, let $\mu_{i}$ and $\nu_{i}$ represent the $s$-vectors $(\mu_{1i},\ldots,\mu_{si})$ and $(\nu_{1i},\ldots,\nu_{si})$, respectively.
For an arbitrary $s$-vector $v=(v_{1},\ldots,v_{s})$ and a real number $\lambda\in\mathbb{R}$, we denote by $\lambda+v$ the $s$-vector $(\lambda+v_{1},\ldots,\lambda+v_{s})$. Given two arbitrary $s$-vectors $v=(v_{1},\ldots,v_{s})$ and $w=(w_{1},\ldots,w_{s})$, $v\circ w$ represents the componentwise product of two $s$-vectors defined by $(v_{1}w_{1},\ldots,v_{s}w_{s})$. We also use the notation $\mathrm{sum}(v)=v_{1}+\cdots+v_{s}$ and $\|v\|_{\infty}=\max(|v_{1}|,\ldots,|v_{s}|)$.
For each $j\in\{1,\ldots,D\}$, we denote by $Y_{n,i}^{j}$, $L_{n,i}^{j}$, and $f^{j}(t_{n-1}+hc_{i},Y_{n,i})$ the $j$th component of $Y_{n,i}\in\mathbb{R}^{D}$, $L_{n,i}\in\mathbb{R}^{D}$ and $f(t_{n-1}+hc_{i},Y_{n,i})\in\mathbb{R}^{D}$ respectively. For each $j\in\{1,\ldots,D\}$, we consider the $s$-vectors $Y_{n}^{j}=(Y_{n,1}^{j},\ldots,Y_{n,s}^{j})$ and $L_{n}^{j}=(L_{n,1}^{j},\ldots,L_{n,s}^{j})$.
Let $Y_{n}$ represent the one-dimensional array $(Y_{n}^{1},\ldots,Y_{n}^{D})$ of $s$-vectors $Y_{n}^{j}$. We define the function $F$ that given $t_{n-1}+hc$ and $Y_{n}$, returns the one-dimensional array $(F_{n}^{1},\ldots,F_{n}^{D})$ of $s$-vectors
$F_{n}^{j}=(f^{j}(t_{n-1}+hc_{1},Y_{n,1}),\ldots,f^{j}(t_{n-1}+hc_{s},Y_{n,s})).$
With that notation, (7) can be written as follows,
$\begin{split}F_{n}&=F(t_{n-1}+hc,Y_{n}),\\ L_{n}^{j}&=h\,(b\circ F_{n}^{j}),\quad j=1,\ldots,D,\\ Y_{n}^{j}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j},\quad j=1,\ldots,D,\end{split}$ (11)
while (8) can be rewritten as
$y_{n}^{j}=y_{n-1}^{j}+\mathrm{sum}(L_{n}^{j}),\quad j=1,\ldots,D.$ (12)
Vectorized implementation of fixed-point iteration
Following (9)–(10), we initialize the fixed-point iteration by approximating $Y_{n}^{j}$ for $j=1,\ldots,D$ as
$Y_{n}^{j,[0]}=y_{n-1}^{j}+\sum_{i=1}^{s}\nu_{i}L_{n-1,i}^{j}.$ (13)
Then, $Y_{n}^{[k]}=(Y_{n}^{1,[k]},\ldots,Y_{n}^{D,[k]})$ is computed for $k=1,2,\ldots$ as follows:
$\begin{split}F_{n}^{[k]}&=F(t_{n-1}+hc,Y_{n}^{[k-1]}),\\ L_{n}^{j,[k]}&=h\,(b\circ F_{n}^{j,[k]}),\quad j=1,\ldots,D,\\ Y_{n}^{j,[k]}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j,[k]},\quad j=1,\ldots,D.\end{split}$ (14)
Stopping criterion
In [12], it is shown that standard stopping criteria based on prescribed iteration error tolerances result in linear growth of energy error when applied to Hamiltonian systems. To address this issue, a new stopping criterion is proposed in [1]. We now present a simplified version of this criterion that is suitable for our vectorized implementation.
If $Y_{n}^{j,[k]}=Y_{n}^{j,[k-1]}$ for all $j\in\{1,\ldots,D\}$, the iteration should clearly be stopped; as $Y_{n}^{j,[k+1]}=Y_{n}^{j,[k]}$ would hold for all $j\in\{1,\ldots,D\}$. However, since this condition may not always be met, the stopping criterion must be supplemented with an additional condition to detect when successive approximations cease to improve: Define $\Delta_{n}^{j,[k]}=\|Y_{n}^{j,[k]}-Y_{n}^{j,[k-1]}\|_{\infty}$. The iteration will stop after the $k$th iteration if the following condition is met for all $j\in\{1,\ldots,D\}$:
$\Delta_{n}^{j,[k]}=0\quad\mbox{or}\quad\min\left(\Delta_{n}^{j,[1]},\ldots,\Delta_{n}^{j,[k-2]}\right)\leqslant\min\left(\Delta_{n}^{j,[k-1]},\Delta_{n}^{j,[k]}\right).$ (15)
Systems of second Order Differential Equations
We now consider systems of second-order differential equations of the form
$\frac{dq}{dt}=v,\quad\frac{dv}{dt}=g(q,t),\quad q(t_{0})=q_{0},\quad v(t_{0})=v_{0},$ (16)
where $q$ and $v$ are vectors with $d$ components each. Hamiltonian systems with a Hamiltonian function of the form
$H(q,p)=\frac{1}{2}\,p^{T}M^{-1}p+U(q)$ (17)
can be expressed in this form via the linear change of variables $p=Mv$.
System (16) is a specific instance of the general form (1), with $D=2d$, $y=(y^{1},\ldots,y^{2d})$, $q=(y^{1},\ldots,y^{d})$, $v=(y^{d+1},\ldots,y^{2d})$, and $f(t,y)=(v,g(t,q))$. It is known [18, 13], that a partitioned form of fixed-point iteration can generally reduce the number of iterations needed for convergence. Such partitioned fixed-point iteration also admits a vectorized implementation that we now describe.
We initialize the fixed-point iteration as
$Y_{n}^{j,[0]}=y_{n-1}^{j}+\sum_{i=1}^{s}\nu_{i}L_{n-1,i}^{j},\quad j=d+1,\ldots,2d.$ (18)
Then, for $k=1,2,\ldots$,
$\begin{split}L_{n}^{j,[k]}&=h\,(b\circ Y_{n}^{d+j,[k-1]}),\quad j=1,\ldots,d,\\ Y_{n}^{j,[k]}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j,[k]},\quad j=1,\ldots,d,\\ F_{n}^{[k]}&=F(t_{n-1}+hc,Y_{n}^{[k]}),\\ L_{n}^{j,[k]}&=h\,(b\circ F_{n}^{j,[k]}),\quad j=d+1,\ldots,2d,\\ Y_{n}^{j,[k]}&=y_{n-1}^{j}+\sum_{i=1}^{s}\mu_{i}L_{n,i}^{j,[k]},\quad j=d+1,\ldots,2d.\end{split}$ (19)
The process will stop after the $k$th iteration if (15) holds for all $j\in\{1,\ldots,d\}$.
3 Implementation aspects of IRKGL16-SIMD
3.1 SIMD-paralellization
Single Instruction, Multiple Data (SIMD) is a parallelization technique supported by modern CPU cores [17, 9]. SIMD enables the execution of vectorized instructions that apply a single operation simultaneously to multiple data elements, thereby accelerating performance in computationally intensive tasks. Modern CPUs contain specialized registers known as short vectors, typically 256 bits (holding four 64-bit double-precision floating-point numbers) or 512 bits (holding eight 64-bit double-precision floating-point numbers) in size.
In a standard SIMD operation, two input vectors are processed element-wise, applying the same operation to each pair of corresponding elements and producing an output vector. By operating on multiple elements in parallel, SIMD significantly improves performance.
3.2 Julia language and SIMD.jl
Our SIMD-vectorized implementation of IRKGL16 (IRKGL16-SIMD) is written in Julia language and relies on the SIMD.jl package. Julia is a high-level dynamic language that allows programmers to write clear, high-level, generic and abstract code resembling mathematical formulas, yet produces fast, low-level machine code that has traditionally only been generated by static languages [3, 4]. By default, Julia employs just-in-time (JIT) compilation, generating LLVM intermediate code that the LLVM compiler framework then translates into optimized machine code.
In Julia, there are several ways to explicitly apply SIMD vectorization, and the SIMD.jl package [19] provides a convenient solution for this purpose. The SIMD.jl library introduces the parameterized vector type Vec{s, T}, representing vectors of $s$ elements of type T. In our vectorized implementation of $s$-stage IRKGL methods (where $s=8$ in IRKGL16-SIMD), we use vectors of type Vec{s, Float64}, meaning $s$-vectors with elements of type ‘Float64‘ (64-bit double-precision floating-point numbers), or alternatively, vectors of type Vec{s, Float32}. The standard arithmetic and logical operations are designed to be applied element-wise in parallel, producing a SIMD vector as the result. SIMD.jl generates LLVM code that defines vectors of $s$ elements, which the LLVM compiler then translates into optimized SIMD instructions that utilize the SIMD registers available on the target platform.
3.3 VecArray and IRKGL16-SIMD
The vectorized form of fixed-point iteration (13)–(14) (resp., (18)–(19)) for first-order problems (1) (resp., second-order problems (16)) is well-suited for SIMD implementation. SIMD-vectorized implementation is particularly effective for IRKGL16, where $s=8$, as this choice maximizes efficient utilization of SIMD registers on modern hardware; in this case, all additions and multiplications in (13)–(14) and (18)–(19) can be implemented efficiently using SIMD operations with short vectors containing eight 64-bit double-precision floating-point numbers.
In addition, the evaluation of $F(t_{n-1}+ch,Y_{n}^{[k-1]})$ in (14) and (19) (which is equivalent to the evaluation of $f(t_{n-1}+c_{i}h,Y_{n,i}^{[k-1]})$ for $i=1,\ldots,s$) also benefits from SIMD parallelization: In IRKGL16-SIMD, parallelization remains transparent to the user, assuming the right-hand side of the system is defined in terms of arithmetic operations and elementary functions compatible with SIMD.jl. Such seamless SIMD parallelization is handled automatically with the help of SIMD.jl and an additional parametrized abstract type, VecArray{s,T}, that we have implemented to efficiently represent and handle abstract arrays of elements of type Vec{s,T} (where T is either Float64 or Float32). The $s$-vectors $b,c,\mu_{i},\nu_{i}$, $Y_{n}^{j}$, $F_{n}^{j}$, and $L_{n}^{j}$ considered in Subsection 2.4 are represented in IRKGL16-SIMD as vectors of type Vec{s,T}, while $Y_{n}=(Y_{n}^{1},\ldots,Y_{n}^{D})$, $F_{n}=(F_{n}^{1},\ldots,F_{n}^{D})$, and $L_{n}=(L_{n}^{1},\ldots,L_{n}^{D})$, and also $\mu=(\mu_{1},\ldots,\mu_{s})$ and $\nu=(\nu_{1},\ldots,\nu_{s})$, are represented as one-dimensional abstract arrays of type VecArray{s,T}.
Let us consider for instance the VecArray object Yn representing $Y_{n}$:
  • The syntax Yn[j] (or equivalently getindex(Yn,j)) can be used to get the vector of type Vec{s,T} representing $Y_{n}^{j}$.
  • The syntax Yn[j] = V (or equivalently setindex!(Yn,j,V)) can be used to update the $j$th vector of type Vec{s,T} representing $Y_{n}^{j}$. Here, we are assuming that V is a vector of type Vec{s,T}.
  • The data of Yn are internally stored as a 2-dimensional array of type Array{T,2} having $s$ rows and $D$ columns.
  • The $s$ components of Yn[j] are stored in the $j$th column of that 2-dimensional array.
Listing LABEL:code:ex1 shows the section of the IRKGL16-SIMD code where a single vectorized fixed-point iteration (14) is implemented.
1f_ODE!(Fn,Yn,parms,tn+h*c)
2
3for j in 1:D
4 Fnj=Fn[j]
5 Lnj=h*(b*Fnj)
6 dYnj=mu[1]*Lnj[1]
7 for i in 2:s
8 dYnj=dYnj+mu[i]*Lnj[i]
9 end
10 Yn[j]=yn[j]+dYnj
11end
The first line of this code requires some explanation, while the remaining lines in Listing LABEL:code:ex1 directly translate the last two lines of (14). The function f_ODE!(dy, y, parms, t) is a user-defined generic function that takes as inputs the array y, a vector parms of constant parameters, and the time t. It computes the $D$ components of the right-hand side $f(t,y)$ of the ODE system, storing the result in the pre-existing array dy. For illustration, Listing LABEL:code:ex2 provides an implementation of f_ODE!(dy, y, parms, t) for the following non-autonomous Hamiltonian ODE system:
$\begin{split}\frac{dq_{1}}{dt}&=p_{1},\quad\frac{dq_{2}}{dt}=p_{2},\\ \frac{dp_{1}}{dt}&=-q_{1}-2(\lambda+\xi\sin(t))q_{1}q_{2},\\ \frac{dp_{2}}{dt}&=-q_{2}-(\lambda+\xi\sin(t))(q_{1}^{2}-q_{2}^{2}).\end{split}$ (20)
Note that this system reduces to the classic Hénon-Heiles system when $\xi=0$. 1function f_ODE!(dy,y,parms,t) 2 3 # q1=y[1]; q2=y[2]; p1=y[3]; p2=y[4] 4 lambda=parms[1] 5 xi=parms[2] 6 aux=lambda+xi*sin(t) 7 dy[1]=y[3] 8 dy[2]=y[4] 9 dy[3]=-y[1]-2*aux*y[1]*y[2] 10 dy[4]=-y[2]-aux*(y[1]^2-y[2]^2) 11 12 return nothing 13end
The generic function f_ODE!(dy, y, parms, t) in Listing LABEL:code:ex2 supports a wide range of argument types. For example, it can be called with t as a scalar of type T = Float64 (or alternatively T=Float32 or T=BigFloat), and dy and y as standard Julia arrays containing elements of type T. Additionally, the function accepts other argument types, provided that the operations within the f_ODE! definition are valid for them. Specifically, the function call f_ODE!(Fn, Yn, parms, tn + h*c) in Listing LABEL:code:ex1, where Fn and Yn are abstract arrays of type VecArray{s, T} (with T = Float64 or T = Float32) and tn + h*c is of type Vec{s, T}, is equivalent to the first line of (14), representing the parallel evaluation of $f(t_{n-1}+c_{i}h,Y_{n,i}^{[k-1]})$ for $i=1,\ldots,s$.
This design enables parallel evaluations of the ODE system’s right-hand side in a way that is transparent to the user, provided that the right-hand side $f(t,y)$ of the ODE system is defined as a generic function f_ODE!(dy, y, parms, t) without branching in terms of arithmetic operations and elementary functions compatible with SIMD.jl. This requirement is met for all the problems considered in the numerical experiments in Section 5 below.
However, not all ODE systems (1) support a generic implementation of f_ODE!(dy, y, parms, t) that enables seamless $s$-fold parallel evaluations of the system’s right-hand side. For instance, this may not be feasible if the definition of $f(t,y)$ involves unavoidable branching, which disrupts the uniformity needed for SIMD processing. In such cases, however, IRKGL16-SIMD can still be applied if the user provides a specialized implementation of f_ODE!(dy, y, parms, t) for instances where dy and y are abstract arrays of type VecArray{s, T} (with T = Float64 or T = Float32) and t is of type Vec{s, T}.
In some cases, such as the $N$-body example considered in Section 5, it is more convenient for the user to organize the state variables in a higher-dimensional array rather than a one-dimensional array. To accommodate this, we have parameterized our abstract array type with an additional parameter dim, as VecArray{s, T, dim}. For one-dimensional arrays of vectors of type Vec{s, T}, we set dim = 2, with the underlying data stored in a two-dimensional standard array of type Array{T, 2}. More generally, VecArray{s, T, dim} refers to $(\texttt{dim}-1)$-dimensional abstract arrays of vectors of type Vec{s, T}, with the data stored in a dim-dimensional array of type Array{T, dim}. Our implementations of the getindex and setindex! functions work as expected in this more general case, allowing users to apply IRKGL16-SIMD with a generic implementation of the function f_ODE!(dy, y, parms, t) when the state variables are stored in a higher-dimensional array.
4 State-of-the-art explicit symplectic integrators
Splitting and composition techniques are powerful tools for constructing practical symplectic integrators in many areas of application. Although the resulting methods are typically tailored to the specific problem and lack the general applicability of approaches such as IRKGL schemes, they can be highly efficient when applicable. [18, 7, 11].
In the context of Hamiltonian systems that can be decomposed into two (or more) exactly solvable components, explicit symplectic integrators can be constructed via operator splitting techniques. A widely used approach is the second-order Strang splitting method [21], which yields a symmetric and symplectic scheme by composing the flows associated with the individual sub-Hamiltonians. Specifically, the Strang splitting scheme advances the solution over a time step $h$ via the composition
$\phi_{h}=\varphi^{A}_{h/2}\circ\varphi^{B}_{h}\circ\varphi^{A}_{h/2}.$ (21)
This method is second-order accurate, time-reversible, and symplectic. The resulting integrator is explicit when the Hamiltonian is written as $H=H_{A}+H_{B}$, and the exact flows $\varphi^{A}_{t}$ and $\varphi^{B}_{t}$ generated by $H_{A}$ and $H_{B}$, respectively, are available in closed form.
For instance, consider a separable Hamiltonian of the form (17), where $H_{A}(q,p)=\tfrac{1}{2}p^{T}M^{-1}p$ and $H_{B}(q,p)=U(q)$. In this case, the exact flows are given by $\varphi^{A}_{t}(q,p)=(q+tM^{-1}p,\,p)$ and $\varphi^{B}_{t}(q,p)=(q,\,p-t\nabla U(q))$.
The method can be extended to systems where the Hamiltonian splits into more than two exactly solvable components. For instance, if $H=H_{A}+H_{B}+H_{C}$ then a symmetric second-order integrator can be obtained by composing the flows of each part as
$\phi_{h}=\varphi^{A}_{h/2}\circ\varphi^{B}_{h/2}\circ\varphi^{C}_{h}\circ\varphi^{B}_{h/2}\circ\varphi^{A}_{h/2}.$ (22)
This generalized Strang-type splitting remains symplectic and time-symmetric, provided that each subflow $\varphi^{A}_{t}$, $\varphi^{B}_{t}$, and $\varphi^{C}_{t}$ is symplectic. Such multi-part splittings arise naturally in applications like charged particle dynamics, Lie-Poisson systems, and systems with coupled fast and slow Hamiltonian components [7, 14].
Extensions to higher-order accuracy can be obtained by symmetric compositions of $s$ basic Strang stages  [7, 11] of the form,
$\psi_{h}=\phi_{\gamma_{s}h}\circ\phi_{\gamma_{s-1}h}\circ\dots\circ\phi_{\gamma_{2}h}\circ\phi_{\gamma_{1}h}$ (23)
with $\gamma_{s+1-j}=\gamma_{j}$, $1\leqslant j\leqslant s$.
We consider the following symmetric compositions of time-symmetric second-order schemes of order $r=2,4,6,8,10$ recommended in [7]:
  • 1)
    Strang splitting: $r=2$th order with $s=1$ stage.
  • 2)
    SUZ90: $r=4$th order with $s=5$ stages proposed by Suzuki(1990) [22].
  • 3)
    SS05: a sixth-order scheme with $s=13$ stages, a eighth-order one with $s=21$ and a tenth-order one with $s=35$ proposed by Sofroniu and Spalleta(2005) [20].
Any composition of the form (23) with (21) can be rewritten as
$\varphi_{a_{s+1}h}^{A}\varphi_{b_{s}h}^{B}\circ\varphi_{a_{s}h}^{A}\circ\dots\circ\varphi_{b_{1}h}^{B}\circ\varphi_{a_{1}h}^{[A]}$ (24)
with coefficients
$a_{1}=\frac{\gamma_{1}}{2},\quad a_{j}=\frac{\gamma_{j-1}+\gamma_{j}}{2},\quad b_{j}=\gamma_{j}.$
Clearly, (24) cannot be written in the form (23) for arbitrary sequences of coefficients $a_{j},b_{j}$. Such more general family of methods are of particular interest in the case of Hamiltonians of the form (17), for which efficient high-order schemes have been constructed. They are often referred as splitting methods of Runge-Kutta-Nyström (RKN) type, as they can be rewritten as explicit symplectic RKN methods applied to the system of second order ODEs (16[18]. We consider the following splitting integrators of RKN type of order $r=6$ and $r=8$ recommended in [7]:
  • 1)
    $\mathrm{S_{ABA}\text{-}BM02}$: a sixth-order method with $s=14$ stages method proposed in Blanes and Moan(2002) [8].
  • 2)
    $\mathrm{S_{ABA}\text{-}BCE22}$: optimized eighth-order with $s=19$ stages proposed in Blanes, Casas and Escorihuela (2022) [6].
5 Numerical experiments
We compare the performance of IRKGL16-SIMD with several state-of-the-art explicit symplectic integrators for the high-precision integration of non-stiff Hamiltonian systems using constant step size. A range of representative examples is considered to assess both accuracy and efficiency.
Some of the explicit symplectic integrators used in the comparison are available through the Julia language’s DifferentialEquations.jl suite [16]. However, for consistency and fair benchmarking, we have implemented all the reference integrators described in the previous section. We verified that, for the problems considered, our implementations of the splitting methods are at least as efficient as—and often more so than—those available in DifferentialEquations.jl.
The source code for the integrators, along with all Jupyter notebooks used to generate the numerical results, is publicly available at: https://github.com/mikelehu/Implicit_symplectic_can_outperform_explicit_symplectic.
All numerical experiments were conducted on an 11th Gen Intel Core i7-11850H processor (2.5 GHz, 16 threads), a CPU equipped with 512-bit SIMD registers. We used the long-term support (LTS) stable release of Julia version 1.10.7 (November 26, 2024) for all tests. To optimize SIMD performance, Julia was launched from the terminal with bounds checking disabled using the flag --check-bounds=no.
5.1 Test problems
We have made comparisons with systems of first-order differential equations as well as with systems of second-order differential equations.
5.1.1 System of first-order differential equations
We consider a Hamiltonian system describing the motion of charged particles in the vicinity of a Schwarzschild black hole under the influence of an external magnetic field, as studied in [15]:
$\begin{split}H(r,\theta,p_{r},p_{\theta})&=\frac{1}{2}\left(1-\frac{2}{r}\right)p_{r}^{2}-\frac{1}{2}\left(1-\frac{2}{r}\right)^{-1}E^{2}+\frac{p_{\theta}^{2}}{2r^{2}}\\ &\quad+\frac{1}{2r^{2}\sin^{2}\theta}\left(L-\frac{\beta}{2}r^{2}\sin^{2}\theta\right)^{2}.\end{split}$ (25)
This Hamiltonian admits multiple splittings into analytically solvable components—typically into three, four, or five parts. Following the findings of [15], we adopt the three-part splitting
$H=H_{A}+H_{B}+H_{C},$
which was shown to offer the best numerical performance across both regular and chaotic trajectories. The individual components are defined as follows:
$H_{A}(r,\theta)$ $=\frac{1}{2r^{2}\sin^{2}\theta}\left(L-\frac{\beta}{2}r^{2}\sin^{2}\theta\right)^{2}-\frac{1}{2}\left(1-\frac{2}{r}\right)^{-1}E^{2},$
$H_{B}(r,p_{r},p_{\theta})$ $=\frac{1}{2}\left(p_{r}^{2}+\frac{p_{\theta}^{2}}{r^{2}}\right),$
$H_{C}(r,p_{r})$ $=-\frac{1}{r}p_{r}^{2}.$
Although [15] provides analytical expressions for the flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$, their direct implementation is computationally inefficient. To enable a fair and representative comparison with our implicit symplectic integrator (IRKGL16), we have improved the evaluation of these flows, significantly enhancing the performance of the explicit method. The derivation and implementation details of this optimized explicit integrator are presented in Appendix A.
As in [15], we set the parameters to $E=0.995$, $L=4.6$, and $\beta=8.9\times 10^{-4}$. The initial conditions are $\theta=\pi/2$ and $p_{r}=0$, with $r=11$ chosen so that the corresponding orbit is regular (i.e., lies on a closed curve). The initial value of $p_{\theta}>0$ is determined by imposing the energy condition $H=-\tfrac{1}{2}$.
5.1.2 Systems of second-order differential equations
We consider two second-order initial value problems of the form (16), corresponding to Hamiltonian systems of the type (17): the outer Solar System model [11] and the Hénon-Heiles problem [12].
6-body outer Solar System.
We examine a simplified model of the outer Solar System, which includes the Sun, the four outer planets (Jupiter, Saturn, Uranus, and Neptune), and Pluto, interacting via mutual Newtonian gravity. The system is described by an 18-degree-of-freedom Hamiltonian, with $q_{i},p_{i}\in\mathbb{R}^{3}$ for $i=1,\dots,6$, and Hamiltonian function
$H(q,p)=\sum_{i=1}^{N}\frac{\|p_{i}\|^{2}}{2m_{i}}-G\sum_{1\leq i<j\leq N}\frac{m_{i}m_{j}}{\|q_{i}-q_{j}\|},$ (26)
where $G$ is the gravitational constant, $m_{i}$ denotes the mass of body $i$, and $N=6$ is the total number of bodies considered.
Initial conditions are taken from the DE430 planetary ephemerides at Julian date 2440400.5 (June 28, 1969) [10], and adjusted so that the barycenter of the system is at rest. The integration is carried out over a time interval of $10^{7}$ days.
Hénon-Heiles system.
We also consider the classical Hénon-Heiles Hamiltonian:
$H(q,p)=\frac{1}{2}(p_{1}^{2}+p_{2}^{2})+\frac{1}{2}(q_{1}^{2}+q_{2}^{2})+q_{1}^{2}q_{2}-\frac{1}{3}q_{2}^{3}.$ (27)
This system is integrated from initial conditions corresponding to regular (non-chaotic) motion over a long time interval of length $2\pi\times 10^{4}$. Specifically, we choose $q_{1}(0)=0$, $q_{2}(0)=0.3$, $p_{2}(0)=0.2$, and determine $p_{1}(0)>0$ such that $H=1/12$ (see [11, 12]).
5.2 Results of Numerical Integrations
5.2.1 Systems of First-Order Differential Equations
Figure 1 presents work-precision diagrams for the Schwarzschild black hole problem. The left panel shows the maximum local error in the Hamiltonian,
$\Delta H^{\text{loc}}_{\max}=\max_{n}\left(\Delta H^{\text{loc}}_{n}\right),\quad\Delta H^{\text{loc}}_{n}=\left|\frac{H(y_{n})-H(y_{n-1})}{H(y_{n-1})}\right|,$ (28)
as a function of the total number of function evaluations. The right panel plots the same error against CPU time. In this plot, the minimal CPU time for the IRKGL16-SIMD method is highlighted with a red vertical dashed line.
We observe that even the sequential implementation of IRKGL16 outperforms the explicit symplectic integrators for the high precision integration of this problem. As for the SIMD implementation, for the same CPU time budget, it achieves significantly higher accuracy even for lower precision integrations.
Figure 1: Work-precision diagrams (maximum local error in the Hamiltonian) for the Schwarzschild black hole problem using the IRKGL16 method and symmetric composition schemes of orders $r=2,4,6,8$, and $10$. In the right plot, the sequential IRKGL16 implementation is denoted as IRKGL16-SEQ
Figure 2 shows the evolution of error over time for different integrators. Each method is integrated using a step size chosen such that the corresponding CPU time approximately matches the red vertical dashed line in the right panel of Figure 1. The left panel displays the global error in the Hamiltonian,
$\Delta H^{\text{glob}}_{\max}=\max_{n}\left(\Delta H^{\text{glob}}_{n}\right),\quad\Delta H^{\text{glob}}_{n}=\left|\frac{H(y_{n})-H(y_{0})}{H(y_{0})}\right|,$ (29)
while the right panel shows the relative error in the radial coordinate:
$\Delta r_{n}=\left|\frac{r_{n}-\tilde{r}_{n}}{\tilde{r}_{n}}\right|,$
where $\tilde{r}_{n}$ denotes the reference solution obtained using IRKGL16-SIMD with a comparatively small step size ($h=16$).
Figure 2: Left: global error in the Hamiltonian. Right: relative error in the radial coordinate $r$ for the Schwarzschild black hole problem. Results are shown for IRKGL16-SIMD and symmetric composition methods of orders $r=2,4,6,8,10$.
5.2.2 Systems of Second-Order Differential Equations
Figure 3 shows work-precision diagrams for the two second-order Hamiltonian problems. The top row corresponds to the 6-body outer Solar System model, while the bottom row pertains to the Hénon-Heiles system. The left panels display the maximum local Hamiltonian error (28) versus function evaluations, and the right panels show the same error versus CPU time. The red dashed line marks the optimal CPU time for the IRKGL16-SIMD method, in the sense that it is the minimal time required to achieve a local accuracy close to round-off error levels.
Figure 3: Work-precision diagrams (maximum local Hamiltonian error) for the 6-body outer Solar System and the Hénon-Heiles problem using IRKGL16, symmetric compositions (orders $r=2,4,6,8,10$), and symplectic RKN methods of orders 6 and 8.
Figures 46 show the evolution of errors, using simulations performed at optimal CPU time. The left panel in Figure 4 corresponds to the global Hamiltonian error of the 6-body outer Solar System, and the right panel to the Hénon-Heiles problem.
Figure 4: Time evolution of the global Hamiltonian error for the 6-body outer Solar System (left) and the Hénon-Heiles problem (right), using IRKGL16, symmetric compositions, and RKN methods.
Figure 5 depicts the position error of each planet in the outer Solar System over time, comparing IRKGL16 and SS05 ($r=10$). Figure 6 shows the error in both position and momentum for the Hénon-Heiles system across the different integrators.
Figure 5: Evolution of position error for each planet in the 6-body outer Solar System using IRKGL16 and SS05 ($r=10$).
Figure 6: Evolution of error in position (left) and momentum (right) for the Hénon-Heiles problem using IRKGL16, symmetric compositions, and RKN methods.
6 Conclusions
IRKGL methods are symplectic and, in non-stiff regimes, can be efficiently implemented using fixed-point iterations. These methods combine symplecticity, time-reversibility, and high-order of accuracy, and they are broadly applicable to general canonical Hamiltonian systems [18]. While optimized high-order explicit symplectic integrators are effective when the Hamiltonian can be decomposed into solvable components, they may be less efficient than IRKGL methods when high-precision integrations are required. The SIMD-based implementation of IRKGL16 significantly enhances performance, often outperforming specialized explicit symplectic methods even at moderate precision levels.
For non-stiff problems with Hamiltonians of the specific form (17), optimized high-order explicit symplectic integrators remain highly effective due to their simplicity and computational efficiency. Although our sequential implementation of IRKGL16 is less efficient than the best available explicit symplectic methods for standard double-precision computations, the SIMD-enhanced version can surpass these methods in performance for high-accuracy integrations, even when constrained to double-precision arithmetic.
[…]
Acknowledgments
All the authors have received funding by the Spanish State Research Agency through project PID2022-136585NB-C22, MCIN/AEI/10.13039/501100011033 and the European Union. They are also partially supported by the Department of Education of the Basque Government through the Consolidated Research Group MATHMODE (ITI456-22).
Appendix A Efficient implementation of the three-part splitting method for the Schwarzschild black hole example
The evaluation of the flows $\varphi_{t}^{A}$ and $\varphi_{t}^{B}$ becomes more efficient when performed in the variables $(x,y,p_{x},p_{y})$, which are related to the original coordinates by the canonical transformation
$x$ $=r\cos\theta,\quad y=r\sin\theta,$
$p_{x}$ $=\cos\theta\,p_{r}-\sin\theta\,\frac{p_{\theta}}{r},\quad p_{y}=\sin\theta\,p_{r}+\cos\theta\,\frac{p_{\theta}}{r}.$
In these new variables, the Hamiltonians $H_{A}$ and $H_{B}$ become
$\bar{H}_{A}(x,y)$ $=\frac{1}{2y^{2}}\left(L-\frac{\beta}{2}y^{2}\right)^{2}-\frac{1}{2}\left(1-\frac{2}{\sqrt{x^{2}+y^{2}}}\right)^{-1}E^{2},$
$\bar{H}_{B}(p_{x},p_{y})$ $=\frac{1}{2}(p_{x}^{2}+p_{y}^{2}).$
For the $H_{C}$ component, the equations of motion are simplified by introducing the variables $(\nu,\mu)$, defined through the non-canonical transformation
$\nu=r\,p_{r},\quad\mu=\frac{p_{r}^{2}}{r}.$
The corresponding equations are
$\frac{d\nu}{dt}=-3\mu,\quad\frac{d\mu}{dt}=0.$
The flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$—and hence the integrators (22) and (23)—can be efficiently evaluated in the variables $(x,y,p_{x},p_{y})$ as follows:
  • 1)
    Flow of $H_{A}$:
    $\varphi_{t}^{A}:(x,y,p_{x},p_{y})\mapsto(x,y,p_{x}^{*},p_{y}^{*}),$
    where
    $p_{x}^{*}$ $=p_{x}-t\,\frac{\partial\bar{H}_{A}}{\partial x}(x,y),$
    $p_{y}^{*}$ $=p_{y}-t\,\frac{\partial\bar{H}_{A}}{\partial y}(x,y).$
  • 2)
    Flow of $H_{B}$:
    $\varphi_{t}^{B}:(x,y,p_{x},p_{y})\mapsto(x^{*},y^{*},p_{x},p_{y}),$
    where
    $x^{*}$ $=x+t\,p_{x},\quad y^{*}=y+t\,p_{y}.$
  • 3)
    Flow of $H_{C}$:
    $\varphi_{t}^{C}:(x,y,p_{x},p_{y})\mapsto(x^{*},y^{*},p_{x}^{*},p_{y}^{*}),$
    where
    $r$ $=\sqrt{x^{2}+y^{2}},\quad c=\frac{x}{r},\quad s=\frac{y}{r},$
    $p_{\theta}$ $=-y\,p_{x}+x\,p_{y},\quad\nu=x\,p_{x}+y\,p_{y},\quad\mu=\frac{\nu^{2}}{r^{3}},$
    $\nu^{*}$ $=\nu-3t\,\mu,\quad p_{r}^{*}=(\mu\,\nu^{*})^{1/3},\quad r^{*}=\frac{\nu^{*}}{p_{r}^{*}},$
    $x^{*}$ $=r^{*}\,c,\quad y^{*}=r^{*}\,s,$
    $p_{x}^{*}$ $=c\,p_{r}^{*}-s\,\frac{p_{\theta}}{r^{*}},\quad p_{y}^{*}=s\,p_{r}^{*}+c\,\frac{p_{\theta}}{r^{*}}.$
The partial derivatives of $\bar{H}_{A}(x,y)$ can be computed efficiently using the intermediate variables:
$z$ $=y^{2},\quad r=\sqrt{x^{2}+z},\quad w=L-\frac{\beta}{2}z,\quad u=1-\frac{2}{r},$
$W$ $=-\frac{w}{z},\quad U=-\frac{E^{2}}{2u^{2}},\quad R=\frac{U}{r^{3}},\quad Z=R+\frac{(W-\beta)W}{2},$
$\frac{\partial\bar{H}_{A}}{\partial x}$ $=-2xR,\quad\frac{\partial\bar{H}_{A}}{\partial y}=-2yZ.$
Unlike the analytical expressions provided in [15] for the flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$, our implementation avoids the use of trigonometric and inverse trigonometric functions entirely, resulting in improved computational efficiency.